IQ scores in a general adult population have mean 100 and standard deviation 15. A simple random sample of 50 adults is selected, and the sample mean x of their IQ scores is computed.
What is the probability that the sample mean IQ score is greater than 105.0? Give your answer as a number between 0 and 1, rounded to 3 decimal places.

Answer by stanbon(75887)

IQ scores in a general adult population have mean 100 and standard deviation 15. A simple random sample of 50 adults is selected, and the sample mean x of their IQ scores is computed.
What is the probability that the sample mean IQ score is greater than 105.0? Give your answer as a number between 0 and 1, rounded to 3 decimal places.
------------
z(105) = (105-100)/15 = 1/3
----
P(x > 105) = P(z > 1/3) = normalcdf(1/3,100) = 0.369
========================================================
Cheers,
Stan H.
==========